OpenStudy (anonymous):

Find f f''(x)=1+3sqrt(x), f(4)=25

OpenStudy (anonymous):

you need one more piece of information to do this

OpenStudy (anonymous):

That's all I was given?

OpenStudy (anonymous):

\[f''(x)=1+3\sqrt{x}\] \[f'(x)=x+2x^{\frac{3}{2}}+c\]

OpenStudy (anonymous):

you can't know \(c\) unless you know some value of \(f'\)

OpenStudy (anonymous):

are you sure it isn't \[f'(x)=1+3\sqrt{x}\] find \(f(x)\)?

OpenStudy (anonymous):

let me check again..

OpenStudy (anonymous):

\[\int\limits_{?}^{?}1+3\sqrt{x}dx = x + 2x ^{3/2}+c\] \[f(4)=25\rightarrow 25= 4+2(4)^{3/2}+c\] \[25=4+16+c \rightarrow c=5\] \[f(x)=x+2x ^{3/2}+5\]